Solve by elimination:
10x+2y=46
3x-3y=21

Answer by mananth(16946)

10 x + 2 y = 46 .............1

3 x -3 y = 21 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 2
30 x 6 y = 138
6 x -6 y = 42
Add the two equations
36 x = 180
/ 36
x = 5
plug value of x in (1)
10 x + 2 y = 46
50 + 2 y = 46
2 y = 46 -50
2 y = -4
y = -2
x= 5, y= -2
m.ananth@hotmail.ca